# Adaptive Data Collection for Latin-American Community-sourced Evaluation of Stereotypes (LACES)

### v1

@@ Abstract
The evaluation of societal biases in NLP models is critically hindered by a glaring geo-cultural gap, as existing benchmarks are overwhelmingly English-centric and focused on U.S. demographics. This leaves regions such as Latin America severely underserved, making it impossible to adequately assess or mitigate the perpetuation of harmful regional stereotypes by language technologies. To address this gap, we introduce a new, large-scale dataset of stereotypes developed through targeted community partnerships across Latin America. Additionally, we present a novel adaptive data collection methodology that uniquely integrates the sourcing of new stereotype entries and the validation of existing data within a single, unified workflow. This combined approach results in a resource with significantly broader coverage and greater regional nuance than static collection methods, enabling a more robust stereotype evaluation and significantly addressing the geo-cultural deficit in fairness resources for Latin America. Furthermore, we believe that our adaptive data collection method could facilitate a more effective, community-engaged approach to capturing a broader spectrum of sociocultural knowledge.
@@
Guido Ivetta1,2, Pietro Palombini1, Sofía Martinelli1, Marcos J. Gomez1,2, Sunipa Dev3, Vinodkumar Prabhakaran3, Luciana Benotti1,2 1Universidad Nacional de Córdoba, Argentina 2Fundación Vía Libre 3Google Research guidoivetta@unc.edu.ar
@@ 1 Introduction
While recent research has begun to address these gaps in evaluation data by curating candidate stereotypes from large language models themselves Bhutani et al. (2024); Jha et al. (2023), this approach anchors primarily on the data and sociocultural knowledge captured in the training set and fails to include more nuanced and broader community insights. In contrast, there is also recent complementary work that engages with a larger set of participants (e.g., Dev et al. (2023); Maina et al. (2024); Mitchell et al. (2025); Ivetta et al. (2025)) to collect a broader set of stereotypes. Our work is in line with these efforts, with a focus on sourcing stereotypes across countries in Latin America.
@@ 2 Related Work
Several benchmark datasets have been proposed to quantify and analyze stereotypes encoded in language models. Early efforts such as CrowS-Pairs (Nangia et al., 2020) and StereoSet (Nadeem et al., 2021) focused on English and U.S.-centric contexts, measuring stereotypical bias across domains like race, religion, gender, and profession. BBQ (Parrish et al., 2022) extended this line of work to question answering, showing that models tend to rely on stereotypes when contextual information is under-specified. More recent resources broaden geographic and linguistic scope. SeeGULL (Jha et al., 2023) used language models to propose stereotypes from different regions and validated them with globally diverse annotators to construct a dataset covering stereotypes from 178 countries across six continents, while SHADES (Mitchell et al., 2025) provided a multilingual parallel dataset spanning 16 languages and 20 regions. Complementary work has emphasized participatory and community-driven approaches: Dev et al. (Dev et al., 2023) engaged with Indian communities to surface locally grounded stereotypes absent from Western-centric benchmarks. Ivetta et al. Ivetta et al. (2025) is a similar effort that engaged Latin American communities. These newer datasets highlight the need for socio-culturally inclusive stereotype evaluation. However, all of them use fixed examples to elicit stereotypes instead of adapting the examples to the social group of the annotator, as we do in this paper.
@@
Different annotators will not necessarily assign the same labels to the same texts, resulting in human label variation Plank (2022). There is evidence that this variation depends on the demographic characteristics of annotators (Binns et al., 2017; Al Kuwatly et al., 2020; Excell and Al Moubayed, 2021; Shen and Rose, 2021). Variation is stronger for subjective tasks like toxic content detection (Sap et al., 2019; Kumar et al., 2021; Sap et al., 2022; Goyal et al., 2022) and stereotype elicitation. Annotation guidelines are known to influence the resulting data obtained. In particular, for concepts such as safety and offensiveness Davani et al. (2024); Aroyo et al. (2023) the definitions of the annotation task are subjective and can be interpreted in different ways. Different definitions of bias and stereotypes have been discussed in previous work, which has found that frequently the definitions are inconsistent or inexistent in work related to these topics in NLP. These definitions are routinely accompanied by examples that illustrate the definition on which annotators rely to understand the definition and from which they generalize. Annotation guidelines play a particularly influential role: the examples included in these documents often serve as prototypes that shape annotators’ interpretations of the task Rogers (2021). Previous work has called for the documentation of data collection guidelines including its examples Bender and Friedman (2018).
@@ 3 Data Collection Methodology
In this section, we first describe the settings in which the data collection was performed, and then we describe the mechanisms.
@@ 3.1 Data Collection Contexts
The data collection was conducted in two settings. First, we organized an in-person workshop at Khipu 2025. We adopted a simultaneous and co-located strategy for data collection, where all participants interacted with the tool in real time. Unlike conventional crowdsourcing, participants were physically present in the same room, performing the task simultaneously. The group included both members of the research community with prior NLP training and newcomers to the field. The workshop lasted 2 hours, facilitated by 3 people with educational training in NLP and ethics. To ensure anonymity while maintaining traceability, each participant was assigned a random identifier not linked to any personal information and was used to label each response. This design encouraged engagement and often provoked surprise, laughter or disagreement reactions with the in-group associations.
Second, the task was also integrated as an activity within a SomosNLP Hackathon, extending the data collection process to an online environment. This allowed us to reach a larger pool of contributors across Latin America and Spain, complementing the in-person workshop, and providing additional validation and new associations.
@@ 3.2 Data Collection Approach
The core task for data collection consisted of validating the given (nationality, attribute) pair as illustrated in Figure 2 and optionally extending the dataset by introducing new pairs, either by proposing a new attribute for the nationality or by associating another nationality to the attribute.
Before starting the annotation process, participants indicated their countries of origin, other countries they felt culturally connected to, and languages they could read and write.
The innovative aspect of the data collection is its design choice to ensure the elicitation of in-group stereotypes. It achieves this by prioritizing presenting annotators with examples drawn from their own country or a neighboring country. A stereotype is in-group if it is generated by somebody from the social group to which the stereotype applies, as defined in Jha et al. (2023). Thus, through this adaptive approach, each annotator is more likely to receive a data point related to their own community.
The annotation interface was designed to be simple and intuitive. At the top of the interface, participants were presented with a (nationality, attribute) pair, where the nationality appeared in red and the attribute in green (see Figure 2). Immediately below, they were asked to evaluate the statement “This is a known association in my region” using a 5-point Likert scale. Participants could decide not to annotate a pair by skipping it.
Participants could provide additional pairs through the optional fields shown in the interface. In (Figure 2, Brazil), they could select other nationalities that they believed were also stereotypically associated to the given attribute. In (Figure 2, make strangers feel like family), participants could propose new attributes for the nationality displayed in the initial pair.
@@
A distinctive feature of our methodology was its adaptive nature. Each new pair generated by participants, was automatically added to the pool of items available for others to evaluate. This mechanism created a continuous feedback loop in which participants not only validated existing associations but also expanded the dataset by observing what others participants added previously. To initialize the task, we relied on a small curated seed set of (nationality, attribute) pairs derived from the HESEIA dataset Ivetta et al. (2025) which focuses in Argentina. The selection of pairs for each participant was random but constrained, introducing prioritization rules to guide the sampling process. In particular, it favored pairs that had received fewer than three validations and, when possible, those related to regions geographically or culturally close to the participant.
The system only presented language pairs the annotator reported understanding. Furthermore, when participants contributed new associations, they were free to do so in any of their declared languages. This setup enabled the construction of a multilingual dataset. Although Spanish dominated due to the demographics of the participant pool, the design allowed for the integration of examples in multiple languages.
While our implementation focused on language and nationality, the same dynamic mechanism could be extended to other participant characteristics (e.g., age or gender) or research objectives, enabling other data collection strategies.
@@
The methodology also aimed to reduce researcher bias by shifting epistemic authority from researchers to annotators. Instead of relying on preselected examples, participants contributed and validated associations dynamically. Disagreements were preserved as meaningful sociolinguistic signals rather than noise. The ultimate goal of LACES is to expose harmful stereotypes in language technologies and address the lack of equity resources for Latin America.
4 Dataset
This dataset covers 120 identities and 842 attributes and it was annotated by 83 individuals from 15 distinct countries. Appendix A shows the specific distribution. These annotators often possessed multiple language proficiencies: 92% claimed to read and write English, apart from their Spanish, or Portuguese native language. The dataset is multilingual and comprises 2437 (nationality, attribute) pairs in Spanish, and 2352 in English.
Our data collection relied on an adaptive methodology, which is why the number of validations is variable across the dataset; they range from just 1 to as many as 63. A subset of 1,262 unique pairs received at least two independent ratings. This is an interesting aspect of the proposed methodology, since we can perform real-time sampling of preferred data points: incentivize geographic proximity, under explored topics, diverse demographic representation, get more validations on controversial topics to better measure agreement, etc.
In this section we will first present the topic categories derived from the annotated attributes with supporting examples. Then, we examine the data validation and resulting inter-annotator disagreement. Finally, we determine which topics demonstrated the strongest consensus and the greatest disagreement.
@@ 4.1 Topics Explored in LACES
To identify the main topics represented in the dataset attributes, a categorization scheme was devised, inspired by the CVQA Romero et al. (2025) and OK-VQA Schwenk et al. (2022) benchmarks. From these existing benchmarks, the categories Cooking and Food, Geography, Buildings, Landmarks, People & Everyday Life, Tradition, Art, History, Sports & Recreation, and Public Figures & Pop Culture were retained. Six additional categories were introduced to better fit the scope of the project: Politics & Governance, Economy, Positive Traits, Negative Traits, Neutral Traits, and Other. Category assignment was initially performed using the GPT-4 OpenAI et al. (2024) model and subsequently reviewed and refined by the research team.
Table 1 shows the set of categories, their frequency, illustrative examples, and the percentage of associations created from in-group (IG) and out-group (OG) perspectives; reflecting whether annotators referred more to their own group or to others. For most topics, associations are generated more frequently about other nationalities, except when referring to local public figures or sports. Economy had the highest percentage of out-group annotations, accounting for 91.12% of its data points.
4.2 Inter-Annotator Disagreement
In this study, the focus was placed on inter-annotator disagreement rather than agreement, examining its relationship with the dataset categories to identify which categories exhibit higher levels of disagreement and may therefore be considered more controversial. Inter-annotator disagreement was quantified using the variance of the scores assigned to each unique (nationality, attribute) association. To assess potential language-specific effects, variance was first computed separately for English and Spanish data; however, no substantial differences were observed and the sample sizes were too small to justify separate reporting, so results are presented jointly.
@@ 4.2 Inter-Annotator Disagreement
For this analysis, only pairs with at least two validations were retained. Variance values ranged from 0 to 4 and showed a highly skewed distribution (skewness = 1.69), with a strong concentration at lower values. Approximately 40% of pairs had zero variance, and the 75th percentile corresponded to a variance below 1. Pairs were classified into two groups: low variance (80%) and high variance (20%). We did not adopt a finer-grained partition since smaller groups would have contained very few samples, complicating subsequent analyses and potentially adding noise into interpretation.
4.3 Consensus and Controversy
@@
Furthermore, we analyzed whether annotator disagreement was associated with specific topics. Relative change for each category between the two groups was computed as $100\times\frac{\text{High}-\text{Low}}{\text{Low}}$, quantifying how much more frequent each category was in the high-variance group. Figure 3 depicts these relative frequency changes. Categories with large positive relative change occurred more frequently in the high-variance group, suggesting greater controversy, whereas categories with negative relative change were more prevalent in the low-variance group, indicating stronger consensus. Notably, Negative Traits emerged as the most controversial category, with a relative change of 182.98%, meaning it was almost three times more prevalent in the high-variance group than in the low-variance group. By contrast, positive and neutral traits did not exhibit comparable patterns of disagreement.
@@
5 Results
This section presents our main findings across three core areas: the novelty of our dataset content, human perception of stereotypes, and the effectiveness of debiasing techniques on language models.
First, we analyze in-group bias in annotator ratings, showing that people agree more when rating positive and neutral attributes associated with their own group. Second, we quantify the dataset conceptual diversity, indicating it contributes a high percentage of unique concepts not covered by similar resources. Finally, we evaluate the performance of zero-shot self-debiasing methods on our data, finding that they are significantly less effective when applied to stereotypes from LACES.
@@ 5 Results
5.1 In-Group Bias and Self-Attribution
The attributes in the LACES dataset were augmented with a sentiment category of Positive, Neutral, or Negative for each attribute through Gemini 2.5 Pro (Comanici et al., 2025). To validate the model’s scoring, a subset of 137 attributes was manually annotated by participants from the same region as the original data. The inter-rater reliability measured by Cohen’s Kappa Cohen (1960) was 0.898, which is typically interpreted as ‘Near-Perfect Agreement’. Figure 4 shows the distribution of scores annotators assigned to stereotypes whose attribute has a Positive, Neutral or Negative sentiment.
This analysis reveals a clear in-group leniency that is mediated by the attribute’s sentiment. When annotators rated stereotypes about their own nationality, the distribution for Positive and Neutral attributes showed a distinct shift toward higher scores (mean = 4.27 and 4.2, respectively), indicating that participants reported these desirable and neutral self-attributions as being significantly better-known in their community than Negative ones (mean = 3.51). The mean recognition score for Positive and Neutral in-group stereotypes is substantially higher than the mean score for any out-group category. Conversely, out-group ratings showed no such sentiment-based variation; recognition scores for out-group stereotypes were uniformly lower across all categories (mean = 3.53, 3.57, 3.43). Participants are more likely to validate the wider social awareness of favorable stereotypes, when those stereotypes apply to their own group.
5.2 Unique Concepts

### v2

@@ Abstract
The evaluation of societal biases in NLP models is critically hindered by a geo-cultural gap, This leaves regions such as Latin America severely underserved, making it impossible to adequately assess or mitigate the perpetuation of harmful regional stereotypes in language technologies.
This paper presents LACES, a stereotype association dataset, for 15 Latin American countries. This dataset includes 4,789 stereotype associations manually created and annotated by 83 participants. The dataset was developed through targeted community partnerships across Latin America.
Additionally, in this paper, we propose a novel adaptive data collection methodology that uniquely integrates the sourcing of new stereotype entries and the validation of existing data within a single, unified workflow. This approach results in a resource with more unique stereotypes than previous static collection methods, enabling a more efficient stereotype collection. The paper further supports the quality of LACES by demonstrating reduced efficacy of debiasing methods on this dataset in comparison to existing popular stereotype benchmarks.
@@
Guido Ivetta1,2, Pietro Palombini1, Sofía Martinelli1, Marcos J. Gomez1,2, M. Emilia Echeveste1,2, Sunipa Dev3, Vinodkumar Prabhakaran3, Luciana Benotti1,2 1Universidad Nacional de Córdoba, Argentina 2Fundación Vía Libre 3Google Research guidoivetta@unc.edu.ar
@@ 1 Introduction
Recent research has begun to address these gaps in evaluation data by curating candidate stereotypes from large language models themselves Bhutani et al. (2024); Jha et al. (2023), although this approach anchors primarily on the data and sociocultural knowledge captured in the training set and fails to include more nuanced and broader community insights. In contrast, there is also recent complementary work that engages with a larger set of participants (e.g., Dev et al. (2023); Maina et al. (2024); Mitchell et al. (2025); Ivetta et al. (2025)) to collect a broader set of stereotypes. Our work is in line with these efforts, with a focus on sourcing stereotypes across countries in Latin America.
@@ 2 Related Work
Several benchmark datasets have been proposed to quantify and analyze stereotypes encoded in language models. Early efforts such as CrowS-Pairs (Nangia et al., 2020) and StereoSet (Nadeem et al., 2021) focused on English and U.S.-centric contexts, measuring stereotypical bias across domains like race, religion, gender, and profession. BBQ (Parrish et al., 2022) extended this line of work to question answering, showing that models tend to rely on stereotypes when contextual information is under-specified. More recent resources broaden geographic and linguistic scope. SeeGULL (Jha et al., 2023) used language models to propose stereotypes from different regions and validated them with globally diverse annotators to construct a dataset covering stereotypes from 178 countries across six continents, while SHADES (Mitchell et al., 2025) provided a multilingual parallel dataset spanning 16 languages and 20 regions. Complementary work has emphasized participatory and community-driven approaches: Dev et al. Dev et al. (2023) engaged with Indian communities to surface locally grounded stereotypes absent from Western-centric benchmarks. Ivetta et al. Ivetta et al. (2025) is a similar effort that engaged Latin American communities. These newer datasets highlight the need for socio-culturally inclusive stereotype evaluation. However, all of them use fixed examples to elicit stereotypes instead of adapting the examples to the social group of the annotator, as we do in this paper.
@@
Different annotators will not necessarily assign the same labels to the same texts, resulting in human label variation Plank (2022a). There is evidence that this variation depends on the demographic characteristics of annotators (Binns et al., 2017; Al Kuwatly et al., 2020; Excell and Al Moubayed, 2021; Shen and Rose, 2021). Variation is stronger for subjective tasks like toxic content detection (Sap et al., 2019; Kumar et al., 2021; Sap et al., 2022; Goyal et al., 2022) and stereotype elicitation. Annotation guidelines are known to influence the resulting data obtained. In particular, for concepts such as safety and offensiveness Davani et al. (2024); Aroyo et al. (2023) the definitions of the annotation task are subjective and can be interpreted in different ways. Different definitions of bias and stereotypes have been discussed in previous work, which has found that frequently the definitions are inconsistent or inexistent in work related to these topics in NLP. These definitions are routinely accompanied by examples that illustrate the definition on which annotators rely to understand the definition and from which they generalize. Annotation guidelines play a particularly influential role: the examples included in these documents often serve as prototypes that shape annotators’ interpretations of the task Rogers (2021). Previous work has called for the documentation of data collection guidelines including its examples Bender and Friedman (2018).
@@ 3 Data Collection Methodology
In this section, we first describe the physical and virtual contexts in which the data was gathered. Then, we detail the data collection task, interactive interface, and adaptive sampling strategies employed to build the dataset.
@@ 3.1 Data Collection Contexts
The data collection was conducted in two settings. First, we organized an in-person workshop at Khipu 2025 (https://khipu.ai/). We adopted a simultaneous and co-located strategy, where all participants interacted with the tool in real time. Unlike conventional crowd-working, participants were physically present in the same room, performing the task simultaneously.
The group included both members of the research community with prior NLP training and newcomers to the field. The workshop lasted 2 hours, facilitated by 3 people with educational training in NLP and ethics. To ensure anonymity while maintaining traceability, each participant was assigned a random identifier not linked to any personal information.
Second, the task was also integrated as an activity within an SomosNLP Hackathon (https://somosnlp.org/), extending the data collection process to an online environment. This allowed us to reach a larger pool of contributors across Latin America and Spain, complementing the in-person workshop, and providing additional validation and new associations.
@@ 3.2 Data Collection Approach
The core task for data collection consisted of validating a given (nationality, attribute) pair, and extending the dataset by introducing new pairs, either by proposing a new attribute for the nationality or by associating another nationality to the attribute.
The annotation interface is illustrated in Figure 2. At the top, participants were presented with a sample pair of (nationality, attribute), where the nationality appeared in red and the attribute in green. Immediately below, they were asked to evaluate the statement “This is a known association in my region” using a 5-point Likert scale. Following this, participants could provide additional pairs through the optional fields shown in the interface. In (Figure 2, Brazil), they could select other nationalities that they believed were also stereotypically associated to the given attribute. In (Figure 2, make strangers feel like family), participants could propose new stereotypical attributes for the nationality displayed in the initial pair.
Our methodology is characterized by its adaptive nature. Each new pair generated by participants was automatically added to the pool of items available for others to evaluate. This mechanism created a continuous feedback loop in which participants not only validated existing associations but also expanded the dataset by observing what other participants added previously.
Before the annotation phase, we collected participant demographics, including country of origin, cultural affiliations, and language proficiency. The selection of input pairs for each participant was governed by a weighted probabilistic sampling algorithm rather than a hard constraint. This approach was designed to balance three primary objectives:
1. In-group Representation: The algorithm weighted geographically or culturally proximate pairs more heavily, ensuring participants primarily evaluated data from their own communities.
2. Validation Coverage: Priority was given to pairs with fewer than three existing validations.
3. Setting-specific Recency: The system prioritized data from the current session to foster real-time feedback and peer engagement.
By employing a probabilistic framework instead of rigid rules, the system maximized in-group validation while maintaining a balanced distribution of annotations, thereby enhancing the overall depth and robustness of the dataset. To initialize the task, we relied on a small and manually curated seed set derived from the HESEIA dataset Ivetta et al. (2025) which focuses in Argentina.
@@
The system only presented pairs in the language the annotator reported understanding, and annotations could be done in any of their declared languages. This setup enabled the construction of a multilingual dataset. Although Spanish dominated due to the demographics of the participant pool, the design allowed for the integration of examples in multiple languages. Furthermore, participants could decide not to annotate a pair by skipping it.
@@
While our implementation focused on language and nationality, the same dynamic mechanism could be extended to other participant characteristics or research objectives, enabling diverse data collection strategies. To facilitate reproducibility and support future research, we have released the source code as an open-source framework. The architecture allows researchers to modify the adaptive sampling logic and the annotation interface to their own data collection needs, providing a versatile infrastructure for various types of community-based, human-in-the-loop initiatives.
4 Dataset Characterization
This dataset covers 120 identities and 842 attributes as annotated by 83 individuals from 15 distinct countries (see Appendix A for nationality distribution). Annotators often possessed multiple language proficiencies: 92% claimed to read and write English, apart from Spanish or Portuguese. The dataset is multilingual and comprises 2437 (nationality, attribute) pairs in Spanish, and 2352 in English.
We employed an adaptive validation methodology, resulting in a dynamic distribution of ratings across the dataset. While all entries attained a minimum of one validation, a subset of 765 unique pairs in Spanish and 501 in English were validated at least twice. This approach allows for real-time sampling and resource optimization, enabling us to prioritize data points based on specific research objectives: incentivizing geographic proximity, capturing under-explored topics, ensuring diverse demographic representation, or increasing the validations for controversial topics to better quantify agreement.
In this section we will first present the explored topics in LACES with supporting examples. Then, we examine the data validation and resulting inter-annotator agreement, determining which topics demonstrated the strongest consensus and the greatest disagreement. Finally, we analyze in-group bias in annotator ratings, showing that people agree more when rating positive and neutral attributes associated with their own group.
@@ 4.1 Topics Explored in LACES
To identify the main topics represented in our dataset, a categorization scheme was devised, inspired by the CVQA Romero et al. (2025) and OK-VQA Schwenk et al. (2022) benchmarks. Six additional categories were introduced to better fit the scope of the project: Politics & Governance, Economy, Positive Traits, Negative Traits, Neutral Traits, and Other. Category assignment was manually reviewed and refined.
Table 1 shows the set of categories, their frequency, illustrative examples, and the percentage of associations created from in-group (IG) and out-group (OG) perspectives; reflecting whether annotators referred more to their own group or to others. Given that 15 countries participated in the study, a purely random sampling approach would yield only 6.67% coverage for in-group annotations. However, since our adaptive methodology specifically incentivized their collection, IG percentages significantly exceed this baseline across all categories.
4.2 Inter-Annotator Agreement
To identify which categories elicit higher levels of disagreement, and may thus be considered more controversial, we quantified inter-annotator agreement by calculating the variance of scores assigned to each unique (nationality, attribute) pair. High variance in these ratings serves as a proxy for sociocultural contention within a given topic.
@@ 4.2 Inter-Annotator Agreement
For this analysis, only pairs with at least two validations were retained. Approximately 40% of pairs had zero variance, and the 75th percentile corresponded to a variance below 1. Pairs were classified into two groups: low variance (80%) and high variance (20%). We did not adopt a finer-grained partition since smaller groups would have contained very few samples, complicating the analysis and potentially adding noise into the interpretation.
@@
We analyzed whether annotator disagreement was associated with specific topics. Relative change for each category between the two groups was computed as $100\times\frac{\text{High}-\text{Low}}{\text{Low}}$, quantifying how much more frequent each category was in the high-variance group.
Figure 3 depicts these relative frequency changes. A large positive change indicates high controversy, whereas a negative change suggests consensus. Notably, Negative Traits emerged as the most controversial category, with a relative change of 182.98%, meaning it was almost three times more prevalent in the high-variance group than in the low-variance group. By contrast, positive and neutral traits did not exhibit comparable patterns of disagreement. On the other side, Sports & Recreation shows the most consensus across topic categories.
@@
4.3 In-Group Bias and Self-Attribution
To further understand how social identity influences the perception of stereotypes, we manually augmented the attributes in the LACES dataset with sentiment labels: Positive, Neutral, or Negative. This allows us to analyze the relationship between an annotator’s group membership and their willingness to validate a stereotype. Figure 4 illustrates the distribution of responses to the prompt: “This is a known association in my region” measured on a 5-point Likert scale (1 = completely unknown, 5 = very well-known).
@@ 4.3 In-Group Bias and Self-Attribution
The results reveal an in-group leniency effect mediated by the sentiment of the attribute. When participants evaluated stereotypes targeting their own nationality (In-Group), we observed that Positive and Neutral attributes showed a distinct shift toward the top of the scale, with mean recognition scores of $4.27$ and $4.20$, respectively. In contrast, participants were notably less likely to acknowledge the social prevalence of Negative stereotypes about their own group, with the mean score dropping to $3.51$. This suggests that individuals are more willing to validate the social prevalence of favorable or benign traits within their own community, likely reflecting a form of collective defensive self-attribution or social desirability bias.
In contrast, the perception of out-group stereotypes showed almost no variation across the three sentiment categories, with means of $3.53$ for Positive, $3.57$ for Neutral, and $3.43$ for Negative attributes. This indicates that while in-group stereotypes are deeply partitioned by the desire to maintain a positive social identity, out-group stereotypes are viewed through a more detached and generalized lens.
By capturing this divergence, LACES offers a unique perspective on the intersection of sentiment and identity in data annotation, highlighting that even well-known stereotypes are filtered through the lens of one’s own community membership.
5 Analysis and Benchmark Comparison
This section compares our dataset with existing benchmarks. First, we quantify the dataset conceptual diversity, indicating it contributes a high percentage of unique concepts not covered by similar resources. Then, we evaluate the performance of self-debiasing methods.
5.1 Unique Attributes